# Age-Gender-Country-Specific Death Rates Modelling and Forecasting: A Linear Mixed-Effects Model

### v1

@@ 5 Exploring the Practical Applications of the LME Model in Life Insurance
Our main goal is to calculate the BEL for both the LME model (87) and the LC model (2) (see Dickson et al., 2019; Wüthrich and Merz, 2013; Gerber, 2013; Promislow, 2014, for an overview). To achieve this, we will fit the models using the Czech mortality dataset, which covers age groups $45$ to $110$ with $1$-year intervals, spanning from $1950$ to $2019$. Subsequently, we will forecast the models for a $120$-year period, enabling us to project the BEL estimates beyond the available data range. To compute the BEL, various factors such as mortality rates, the insured individual’s age, maximum age ($110$ years), retirement age ($65$ in the Czech Republic), valuation year, interest rate, and an indicator for discounting premium or annuity (type) are taken into account. In this analysis, we will utilize a valuation year of $2023$ and a interest rate of $0.1\%$ These inputs will be utilized in the calculations to estimate the BEL for the specified scenario.
@@
Beyond computing the BEL, we also conduct simulations to determine the Solvency Capital Requirement (SCR), which represents the capital amount an insurance company must hold to survive a 1 in 200 event (see Wüthrich and Merz, 2013; Promislow, 2014; Dickson et al., 2019, for an overview). Our analysis produces four barplots. The initial two barplots display the BEL values for the LME model and the LC model, yielding $1,935,583$ CZK and $1,936,782$ CZK, respectively, see Figure 27.
@@ Appendix C A hypothetical portfolio and experience factors
Table 2: Hypothetical experience factors: This table presents a hypothetical Experience Factors table for age groups ranging from 00 to $110$, specifically designed for an insurance company operating in the Czech Republic.

### v2

@@ 5 Exploring the Practical Applications of the LME Model in Life Insurance
Our main goal is to calculate the BEL for both the LME model (23) and the LC model (2) (see dickson2019actuarial; wuthrich2013financial; gerber2013life; promislow2014fundamentals, for an overview). To achieve this, we will fit the models using the Czech mortality dataset, which covers age groups $45$ to $110$ with $1$-year intervals, spanning from $1950$ to $2019$. Subsequently, we will forecast the models for a $120$-year period, enabling us to project the BEL estimates beyond the available data range. To compute the BEL, various factors such as mortality rates, the insured individual’s age, maximum age ($110$ years), retirement age ($65$ in the Czech Republic), valuation year, interest rate, and an indicator for discounting premium or annuity (type) are taken into account. In this analysis, we will utilize a valuation year of $2023$ and a interest rate of $0.1\%$ . These inputs will be utilized in the calculations to estimate the BEL for the specified scenario.
@@
Beyond computing the BEL, we also conduct simulations to determine the Solvency Capital Requirement (SCR), which represents the capital amount an insurance company must hold to survive a 1 in 200 event (see wuthrich2013financial; promislow2014fundamentals; dickson2019actuarial, for an overview). Our analysis produces four barplots. The initial two barplots display the BEL values for the LME model and the LC model, yielding $1,935,583$ CZK and $1,936,782$ CZK, respectively, see Figure 27.
@@ Appendix C A hypothetical portfolio and experience factors
Table 2: Hypothetical experience factors: This table presents a hypothetical Experience Factors table for age groups ranging from 0 to $110$, specifically designed for an insurance company operating in the Czech Republic.